40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5632     Entry time: Fri Oct 7 19:06:46 2011
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC spot positions: checked and corrected. 

Koji and Kiwamu had adjusted the MC beam axis slightly such that we can couple the MC output into the Y-arm without exceeding the current range of adjustability on PZT1.  This changed the centering of beam spots on MC mirrors.  I checked the mc-decentering make sure we have not made too big a compromise.  And since we can move MC2 spot position while maintaining the current positions on MC1 and MC3 decentering, we can atleast eliminate the A2L coupling on that mirror.  I used the scripts in $scripts$/MC/moveMC2/ to adjust the MC2 spot position.

Spot positions in mm (MC1,2,3 pit MC1,2,3 yaw) before adjustment:
    1.4674   -0.3548    1.0199   -1.5519    1.9834   -1.5971

After correcting MC2:

    1.4528    0.1431    0.9958   -1.2147    0.3823   -2.0163

After correcting MC1:

    1.3745    0.0669    0.8899   -1.5269    0.0296   -1.7314

 

The spot positions on MC1 and MC3 are very nearly (+/- 0.06 mm) same as before, while the MC2  decentering has been reduced close to zero.

A slight adjustment of the PZTs may be required to reset the beam pointing.

ELOG V3.1.3-